# Next-Generation Database Interfaces: A Survey of LLM-based Text-to-SQL

### v5

@@ Abstract
Generating accurate SQL from users’ natural language questions (text-to-SQL) remains a long-standing challenge due to the complexities involved in user question understanding, database schema comprehension, and SQL generation. Traditional text-to-SQL systems, which combine human engineering and deep neural networks, have made significant progress. Subsequently, pre-trained language models (PLMs) have been developed for text-to-SQL tasks, achieving promising results. However, as modern databases and user questions grow more complex, PLMs with a limited parameter size often produce incorrect SQL. This necessitates more sophisticated and tailored optimization methods, which restricts the application of PLM-based systems. Recently, large language models (LLMs) have shown significant capabilities in natural language understanding as model scale increases. Thus, integrating LLM-based solutions can bring unique opportunities, improvements, and solutions to text-to-SQL research. In this survey, we provide a comprehensive review of existing LLM-based text-to-SQL studies. Specifically, we offer a brief overview of the technical challenges and evolutionary process of text-to-SQL. Next, we introduce the datasets and metrics designed to evaluate text-to-SQL systems. Subsequently, we present a systematic analysis of recent advances in LLM-based text-to-SQL. Finally, we make a summarization and discuss the remaining challenges in this field and suggest expectations for future research directions.
@@ I Introduction
Text-to-SQL is a long-standing task in natural language processing, aiming to convert natural language (NL) questions into database-executable SQL queries. Fig. 1 illustrates an example of a large language model-based (LLM-based) text-to-SQL system. Given a user question such as “Could you tell me the names of the 5 leagues with the highest matches of all time and how many matches were played in the said league?”, the LLM takes the question and its corresponding database schema as input to generate an SQL query. Executing this SQL query in the database retrieves the relevant content needed to answer the user’s question. This system serves as a natural language interface to the database (NLIDB) using LLMs. Since SQL is one of the most widely used programming languages, with over half (51.52%) of professional developers employing it, but only around a third (35.29%) are systematically trained [^1] https://survey.stackoverflow.co/2023, the NLIDB allows non-skilled users to access structured databases like professional database engineers [2, 3] and enhances human-computer interaction [4]. Furthermore, amid the research hotspot of LLMs, text-to-SQL can potentially mitigate prevalent hallucination issues by integrating realistic content from databases to fill LLMs’ knowledge gaps [5]. The significant value of text-to-SQL have triggered a range of studies on its integration and optimization with LLMs [6, 7, 8]; consequently, LLM-based text-to-SQL remains a highly discussed research field in the NLP and database communities.
@@
The overall implementation details of LLM-based text-to-SQL can be divided into three aspects: 1. Question understanding: The NL question is a semantic representation of the user’s intention, which the corresponding generated SQL query is expected to align with; 2. Schema comprehension: The schema provides the table and column structure of the database, and the text-to-SQL system is required to identify the target components that match the user question; 3. SQL generation: This involves incorporating the above parsing and then predicting the correct syntax to generate executable SQL queries that can retrieve the required answer. LLMs have proven to perform a successful vanilla implementation [6, 25], benefiting from the powerful semantic parsing capacity enabled by rich training corpus [26, 27]. Further studies on enhancing LLMs for question understanding [7, 8], schema comprehension [28, 29], and SQL generation [30] are being increasingly released.
@@
CHESS [56], SuperSQL [57], Distillery [58], RSL-SQL [59], Gen-SQL [60], ReFoRCE [61], modelnode-methods,text width=42.5em] ] [ Prompt
@@
SAFE-SQL [72], ReFoRCE [61], modelnode-methods,text width=42.5em] ] [ Reasoning
@@
SQL-CRAFT [74], FUXI [75], SuperSQL [57], CHASE-SQL [76], ReFoRCE [61], modelnode-methods,text width=42.5em] ] [ Execution
@@
MCS-SQL [55], CHESS [56], Distillery [58], E-SQL [80], CHASE-SQL [76], RSL-SQL [59], modelnode-methods,text width=42.5em] ] ] [ Fine-tuning
@@
XiYan-SQL [84], modelnode-methods,text width=42.5em] ] [ Multi-task
Tuning, leaf-methods, text width=6.5em [SQL-LLaMA [85], DTS-SQL [86], KaSLA [29], MSc-SQL [87], XiYan-SQL [84], ROUTE [88], modelnode-methods,text width=42.5em] ] ] ] ]
@@
To summarize our technical contributions, at the time of this paper’s release, we present the first LLM-based text-to-SQL survey in the community, establishing a systematic taxonomy for LLM-based text-to-SQL methods. This categorization provides analysis from technical implementation perspectives, clearly indicating their contributions to generating better SQL. Then, we further integrate insights while analyzing the trade-offs between LLM-based and traditional methods, as well as between in-context learning and fine-tuning paradigms for future developments. Finally, we highlight the remaining challenges in robustness, real-world deployment, efficiency, and domain adaptation, offering actionable directions for future research. Fig. 3 shows a taxonomy tree that summarizes the structure and contents of our survey.
@@ II Overview
Text-to-SQL is a task that aims to convert NL questions into SQL queries that can be executed in a relational database. It has the potential to democratize data access by allowing users to interact with databases using natural language, thus eliminating the need for specialized SQL programming knowledge [93]. This capability can greatly benefit domains such as business intelligence, customer support, and scientific research, by enabling non-specialized users to easily retrieve target content from databases and facilitating more efficient data analysis.
@@ III-A2 Knowledge-augmented Dataset
Interest in incorporating domain-specific knowledge into text-to-SQL tasks has increased significantly in recent years. BIRD [1] employs human database experts to annotate each text-to-SQL sample with external knowledge, categorized into numeric reasoning knowledge, domain knowledge, synonym knowledge, and value illustration. Similarly, Spider-DK [36] defines and adds five types of domain knowledge for a human-curated version of the Spider dataset [11]: SELECT columns mentioned by omission, simple inference required, synonyms substitution in cell value word, one non-cell value word generate a condition, and easy to conflict with other domains. Both studies found that human-annotated knowledge significantly improves SQL generation performance for samples requiring external domain knowledge. Additionally, SQUALL [41] manually annotates alignments between the words in NL questions and the entities in SQL, providing finer-grained supervision than other datasets.
@@ III-A3 Context-dependent Dataset
SParC [40] and CoSQL [32] explore context-dependent SQL generation by constructing a conversational database querying system. Unlike traditional text-to-SQL datasets that only have a single question-SQL pair for one example, SParC decomposes the question-SQL examples in the Spider dataset into multiple sub-question-SQL pairs to construct a simulated and meaningful interaction, including inter-related sub-questions that aid SQL generation, and unrelated sub-questions that enhance data diversity. CoSQL, in comparison, involves conversational interactions in natural language, simulating real-world scenarios to increase complexity and diversity. Additionally, Spider-SS&CG [35] splits the NL question in the Spider dataset [11] into multiple sub-questions and sub-SQLs, demonstrating that training on these sub-examples can improve a text-to-SQL system’s generalization ability on out-of-distribution samples.
@@ III-A4 Robustness Dataset
Evaluating the accuracy of text-to-SQL systems with polluted or perturbed database contents (e.g., schema and tables) is crucial for assessing robustness. Spider-Realistic [38] removes explicit schema-related words from the NL questions, while Spider-SYN [37] replaces them with manually selected synonyms. ADVETA [34] introduces adversarial table perturbation (ATP), which perturbs tables by replacing original column names with misleading alternatives and inserting new columns with high semantic associations but low semantic equivalency. Based on the generation ability of large-scale pre-trained language models, Dr. Spider [44] provided 17 different perturbation strategies on databases, natural language queries and SQL statements to measure the robustness of text-to-SQL system. These perturbations lead to significant drops in accuracy, as a text-to-SQL system with low robustness may be misled by incorrect matches between tokens in NL questions and database entities.
@@ III-A6 Long-context dataset
In real-world database applications, numerous challenging scenarios demand complex SQL queries with significant token lengths, often exceeding 100 tokens, involving multiple intricate operations. Spider 2.0 [42] exemplifies this complexity by presenting enterprise-level text-to-SQL problems that require advanced reasoning across various SQL queries and dialects. This benchmark reveals the limitations of current models, as demonstrated by lower success rates—with advanced LLMs solving only 17.1% of tasks—underscoring its unique challenges. In parallel, BIRD-CRITIC [^2] https://bird-critic.github.io/ [1] introduces a diagnostic dimension, comprising 600 tasks designed to assess the efficacy of LLMs in resolving real-world user issues across diverse SQL dialects. Together, these datasets push the boundaries of text-to-SQL research, emphasizing the challenges of complex, long-context SQL environments.
@@ III-A7 Specialized-domain dataset
Despite advancements in text-to-SQL research, there is still a shortage of specialized datasets tailored for domain-specific analysis. To address this gap, BULL [45] introduces a dedicated text-to-SQL benchmark for the financial sector. The dataset encompasses databases related to funds, stocks, and macroeconomic data, offering a comprehensive platform to evaluate and enhance text-to-SQL methods in financial applications, effectively tackling the financial domain unique text-to-SQL challenges.
@@ III-B2 Execution-based Metrics
    Then, $\textbf{R}(Y_{n},\hat{Y}_{n})=\sqrt{E(Y_{n})/E(\hat{Y}_{n})}$ denotes the relative execution efficiency of the predicted SQL query in comparison to ground-truth query, where $E(\cdot)$ is the execution time of each SQL in the database. BIRD benchmark [1] ensures the stability of this metric by computing the average of $\textbf{R}(Y_{n},\hat{Y}_{n})$ over 100 runs for each example.
Most of the recent LLM-based text-to-SQL studies focus on BIRD and Spider-related datasets and execution-based metrics, the corresponding details can be found at Tab. III.
@@ IV-A In-context Learning
Through extensive and widely recognized research, ICL (also called prompt engineering) has been proven to play a decisive role in the performance of LLMs across various tasks [117, 26, 118]. Consequently, it also impacts SQL generation across different prompt styles [8, 119]. The ICL implementation of the LLM-based text-to-SQL process to generate a predicted SQL query $\hat{Y}$, can be formulated as:
@@
where $k$ is the number of provided instances (examples) in few-shot prompting. $\mathcal{E}_{i}$ represents the $i$-th few-shot instance, each of which can be decomposed as $\mathcal{E}_{i}=\left(\mathcal{S}_{i},Q_{i},Y_{i}\right)$, where $Y_{i}$ is the gold SQL query of $Q_{i}$. The index $i\in\{1,\ldots,k\}$ also serves as the corresponding serial number for $\mathcal{S}_{i}$, $Q_{i}$, and $Y_{i}$. The OpenAI demonstration [^3] The prompt format follows the official document provided by the OpenAI platform: https://platform.openai.com/examples/default-sql-translate provides an indicative example and regulation for the prompting process in LLM-based text-to-SQL [28]. The prompt designs and formats utilized by a substantial portion of the ICL methods discussed in this survey are developed based on the OpenAI regulation. Furthermore, most of these ICL methods incorporate zero-shot and few-shot prompting into their text-to-SQL frameworks.
@@ IV-A1 $\mathbf{C}_{0}$-Vanilla Prompting
TABLE II: A summary of representative methods used in the ICL paradigm for LLM-based text-to-SQL is provided. The complete table, along with a detailed categorization of existing well-designed methods denoted as $\mathbf{C}_{1:4}$, is available in Tab. III. *Note that SimpleDDL is a prompting style that provides only table and column names [119], rather than a method.
@@
TABLE III: Well-designed methods in the ICL paradigm for LLM-based text-to-SQL are ordered by release time. The methods are grouped into four categories based on their implementation perspective: $\mathbf{C}_{1}$-Decomposition, $\mathbf{C}_{2}$-Prompt Optimization, $\mathbf{C}_{3}$-Reasoning Enhancement, and $\mathbf{C}_{4}$-Execution Refinement. Methods that fall into multiple categories will be introduced accordingly. The dataset column consists of six text-to-SQL datasets, which can be clarified by names: Spider [11], BIRD [1], Spider-DK [36] Spider-SYN [37], Spider-Realistic [38], and Spider 2.0 [42]. We use the reference ID in the table due to column space limitations. *There are multiple applied LLMs in the corresponding method; the model with representative performance is presented. CoT method is introduced in multiple venues: NeurIPS’23 [1], EMNLP’23 [47], VLDB’24 [8], arXiv’24 [30].
Methods Applied LLMs Dataset Metrics $\mathbf{C}_{1}$ $\mathbf{C}_{2}$ $\mathbf{C}_{3}$ $\mathbf{C}_{4}$ Release Time Publication Venue
MBR-Exec [77] Codex [11] EX Apr-2022 EMNLP’22
Coder-Reviewer [46] Codex [11] EX Nov-2022 ICML’23
LEVER [78] Codex [11] EX Feb-2023 ICML’23
Self-Debugging [79] StarCoder* [11] EX Apr-2023 ICLR’24
DESEM+P [62] ChatGPT [11, 37] EX Apr-2023 PRICAI’23
DIN-SQL [7] GPT-4* [1, 11, 42] EX, EM, VES Apr-2023 NeurIPS’23
CoT [47, 1, 8, 30] GPT-4 [11, 38, 1] EX, VES May-2023 Multiple Venues
StructGPT [63] ChatGPT* [11, 37, 38] EX May-2023 EMNLP’23
SD+SA+Voting [64] ChatGPT* [11, 37, 38] EX May-2023 EMNLP’23 Findings
QDecomp [47] Codex [11, 38] EX May-2023 EMNLP’23
Least-to-Most [47] Codex [11] EX May-2023 EMNLP’23
SQL-PaLM [73] PaLM-2 [11] EX May-2023 TMLR’24
RAG+SP&DRC [65] ChatGPT [11] EX Jul-2023 ICONIP’23
C3 [28] ChatGPT [11] EX Jul-2023 arXiv’23
DAIL-SQL [8] GPT-4* [11, 1, 38, 42] EX, EM, VES Aug-2023 VLDB’24
ODIS [66] Codex* [11] EX Oct-2023 EMNLP’23 Findings
ACT-SQL [67] GPT-4* [11, 37] EX, EM Oct-2023 EMNLP’23 Findings
MAC-SQL [48] GPT-4* [11, 1] EX, EM, VES Dec-2023 COLING’25
DEA-SQL [49] GPT-4 [11] EX Feb-2024 ACL’24 Findings
FUSED [68] ChatGPT* [11] EX Feb-2024 EMNLP’24 Findings
DELLM [69] GPT-4* [11, 1] EX, VES Feb-2024 ACL’24 Findings
SGU-SQL [30] GPT-4* [11, 1] EX, EM Feb-2024 arXiv’24
PoT [74] GPT-4* [11, 1] EX Feb-2024 arXiv’24
SQL-CRAFT [74] GPT-4* [11, 1] EX Feb-2024 arXiv’24
FUXI [75] GPT-4* [1] EX Feb-2024 EMNLP’24
MetaSQL [50] GPT-4* [11] EX, EM Feb-2024 ICDE’24
$\textit{{R}}^{3}$ [51] GPT-4* [11, 1] EX Feb-2024 arXiv’24
PET-SQL [52] GPT-4 [11] EX Mar-2024 arXiv’24
PURPLE [53] GPT-4* [11, 37, 38] EX, EM Mar-2024 ICDE’24
Dubo-SQL [70] ChatGPT* [1] EX Apr-2024 arXiv’24
TA-SQL [70] GPT-4* [11, 1, 36, 38] EX Apr-2024 ACL’24 Findings
MCS-SQL [55] GPT-4 [11, 1] EX, VES May-2024 COLING’25
CHESS [56] GPT-4o* [11, 1, 42] EX, VES May-2024 arXiv’24
SuperSQL [57] GPT-4* [11, 1] EX, EM, VES Jun-2024 VLDB’24
Distillery [58] GPT-4o-Mini* [1] EX Aug-2024 NeurIP’24 Workshop
E-SQL [80] GPT-4o* [11, 1] EX, VES Sep-2024 arXiv’24
CHASE-SQL [76] Gemini-1.5* [11, 1] EX, VES Oct-2024 ICLR’25
RSL-SQL [59] DeepSeek-Chat* [11, 1] EX, VES Nov-2024 arXiv’24
ICRL [71] ChatGPT [11] EX Jan-2025 COLING’25
Gen-SQL [60] LLaMA-3* [11, 1] EX, EM, VES Jan-2025 COLING’25
SAFE-SQL [72] GPT-4o-Mini* [11] EX, EM Feb-2025 arXiv’25
ReFoRCE [61] o1-preview* [42] EX Feb-2025 arXiv’25

### v6

@@ Abstract
Generating accurate SQL from users’ natural language questions (text-to-SQL) remains a long-standing challenge due to the complexities involved in user question understanding, database schema comprehension, and SQL generation. Traditional text-to-SQL systems, which combine human engineering and deep neural networks, have made significant progress. Subsequently, pre-trained language models (PLMs) have been developed for text-to-SQL tasks, achieving promising results. However, as modern databases and user questions grow more complex, PLMs with a limited parameter size often produce incorrect SQL. This necessitates more sophisticated and tailored optimization methods, which restrict the application of PLM-based systems. Recently, large language models (LLMs) have shown significant capabilities in natural language understanding as model scale increases. Thus, integrating LLM-based solutions can bring unique opportunities, improvements, and solutions to text-to-SQL research. In this survey, we provide a comprehensive review of existing LLM-based text-to-SQL studies. Specifically, we offer a brief overview of the technical challenges and evolutionary process of text-to-SQL. Next, we introduce the datasets and metrics designed to evaluate text-to-SQL systems. Subsequently, we present a systematic analysis of recent advances in LLM-based text-to-SQL. Finally, we make a summary and discuss the remaining challenges in this field and suggest expectations for future research directions. All the related resources of LLM-based, including research papers, benchmarks, and open-source projects, are collected for the community in our repository: https://github.com/DEEP-PolyU/Awesome-LLM-based-Text2SQL.
@@ I Introduction
Text-to-SQL is a long-standing task in natural language processing, aiming to convert natural language (NL) questions into database-executable SQL queries. Fig. 1 illustrates an example of a large language model-based (LLM-based) text-to-SQL system. Given a user question such as “Could you tell me the names of the 5 leagues with the highest matches of all time and how many matches were played in the said league?”, the LLM takes the question and its corresponding database schema as input to generate an SQL query. Executing this SQL query in the database retrieves the relevant content needed to answer the user’s question. This system serves as a natural language interface to the database (NLIDB) using LLMs. According to the 2023 Stack Overflow survey, SQL is one of the most widely used programming languages, with over half of professional developers (51.52%) employing it, while only about a third (35.29%) having received systematic training. The NLIDB therefore allows non-skilled users to access structured databases like professional database engineers [2, 3] and enhances human-computer interaction [4]. Furthermore, amid the research hotspot of LLMs, text-to-SQL can potentially mitigate prevalent hallucination issues by integrating realistic content from databases to fill LLMs’ knowledge gaps [5]. The significant value of text-to-SQL has triggered a range of studies on its integration and optimization with LLMs [6, 7, 8]; consequently, LLM-based text-to-SQL remains a highly discussed research field in the NLP and database communities.
@@
The overall implementation details of LLM-based text-to-SQL can be divided into three aspects: 1. Question understanding: The NL question is a semantic representation of the user’s intention, which the corresponding generated SQL query is expected to align with; 2. Schema comprehension: The schema provides the table and column structure of the database, and the text-to-SQL system is required to identify the target components that match the user question; 3. SQL generation: This involves incorporating the above parsing and then predicting the correct syntax to generate executable SQL queries that can retrieve the required answer. LLMs have proven to perform a successful vanilla implementation [6, 25], benefiting from the powerful semantic parsing capacity enabled by a rich training corpus [26, 27]. Further studies on enhancing LLMs for question understanding [7, 8], schema comprehension [28, 29], and SQL generation [30] are being increasingly released.
@@
CHESS [56], SuperSQL [57], Distillery [58], RSL-SQL [59], Gen-SQL [60], Spider-Agent [42],
ReFoRCE [61], LinkAlign [62], modelnode-methods,text width=42.5em] ] [ Prompt
@@
SAFE-SQL [73], ReFoRCE [61], LinkAlign [62], modelnode-methods,text width=42.5em] ] [ Reasoning
@@
SQL-CRAFT [75], FUXI [76], SuperSQL [57], CHASE-SQL [77], ReFoRCE [61], Spider-Agent [42], modelnode-methods,text width=42.5em] ] [ Execution
@@
MCS-SQL [55], CHESS [56], Distillery [58], E-SQL [81], CHASE-SQL [77], RSL-SQL [59],
Spider-Agent [42], ReFoRCE [61], modelnode-methods,text width=42.5em] ] ] [ Fine-tuning
@@
XiYan-SQL [85], SHARE [86], modelnode-methods,text width=42.5em] ] [ Multi-task
Tuning, leaf-methods, text width=6.5em [SQL-LLaMA [87], DTS-SQL [88], KaSLA [29], MSc-SQL [89], XiYan-SQL [85], ROUTE [90],
SHARE [86], modelnode-methods,text width=42.5em] ] ] ] ]
@@
To summarize our technical contributions, at the time of this paper’s release, we present the first LLM-based text-to-SQL survey in the community, establishing a systematic taxonomy for LLM-based text-to-SQL methods. This categorization provides analysis from technical implementation perspectives, clearly indicating their contributions to generating better SQL. Then, we further integrate insights while analyzing the trade-offs between LLM-based and traditional methods, as well as between in-context learning and fine-tuning paradigms for future research. Finally, we highlight the remaining challenges in robustness, real-world deployment, efficiency, and domain adaptation, offering actionable directions for future research. Fig. 3 shows a taxonomy for the survey.
@@ II Overview
Text-to-SQL is a task that aims to convert NL questions into SQL queries that can be executed in a relational database. It has the potential to democratize data access by allowing users to interact with databases using natural language, thus eliminating the need for specialized SQL programming knowledge [95]. This capability can greatly benefit domains such as business intelligence, customer support, and scientific research by enabling non-specialized users to easily retrieve target content from databases and facilitating more efficient data analysis.
@@ III-A2 Knowledge-augmented Dataset
Interest in incorporating domain-specific knowledge into text-to-SQL tasks has increased significantly in recent years. BIRD [1] employs human database experts to annotate each text-to-SQL sample with external knowledge, categorized into numeric reasoning knowledge, domain knowledge, synonym knowledge, and value illustration. Similarly, Spider-DK [36] defines and adds five types of domain knowledge for a human-curated version of the Spider dataset [11]: SELECT columns mentioned by omission, simple inference required, synonym substitution in cell value words, one non-cell value word generates a condition, and easy to conflict with other domains. Both studies found that human-annotated knowledge significantly improves SQL generation performance for samples requiring external domain knowledge. Additionally, SQUALL [41] manually annotates alignments between the words in NL questions and the entities in SQL, providing finer-grained supervision than other datasets.
@@ III-A3 Context-dependent Dataset
SParC [40] and CoSQL [32] explore context-dependent SQL generation by constructing a conversational database querying system. Unlike traditional text-to-SQL datasets that only have a single question-SQL pair for one example, SParC decomposes the question-SQL examples in the Spider dataset into multiple sub-question-SQL pairs to construct a simulated and meaningful interaction, including interrelated sub-questions that aid SQL generation, and unrelated sub-questions that enhance data diversity. CoSQL, in comparison, involves conversational interactions in natural language, simulating real-world scenarios to increase complexity and diversity. Additionally, Spider-SS&CG [35] splits the NL question in the Spider dataset [11] into multiple sub-questions and sub-SQLs, demonstrating that training on these sub-examples can improve a text-to-SQL system’s generalization ability on out-of-distribution samples.
@@ III-A4 Robustness Dataset
Evaluating the accuracy of text-to-SQL systems with polluted or perturbed database contents (e.g., schema and tables) is crucial for assessing robustness. Spider-Realistic [38] removes explicit schema-related words from the NL questions, while Spider-SYN [37] replaces them with manually selected synonyms. ADVETA [34] introduces adversarial table perturbation (ATP), which perturbs tables by replacing original column names with misleading alternatives and inserting new columns with high semantic associations but low semantic equivalency. Based on the generation ability of large-scale pre-trained language models, Dr. Spider [44] provided 17 different perturbation strategies on databases, natural language queries and SQL statements to measure the robustness of text-to-SQL systems. These perturbations lead to significant drops in accuracy, as a text-to-SQL system with low robustness may be misled by incorrect matches between tokens in NL questions and database entities.
@@ III-A6 Long-context dataset
In real-world database applications, numerous challenging scenarios demand complex SQL queries with significant token lengths, often exceeding 100 tokens and involving multiple intricate operations. Spider 2.0 [42] exemplifies this complexity by presenting enterprise-level text-to-SQL problems that require advanced reasoning across various SQL queries and dialects. Distinct from earlier benchmarks, Spider 2.0 contains numerous real-world tasks sourced from diverse production databases such as BigQuery and Snowflake, featuring large-scale, industrial schemas with hundreds or even thousands of columns, nested structures, and multiple schemas per instance. Each task requires navigating not only complex SQL logic, but also integrating information from project codebases, database metadata, and dialect-specific documentation, closely reflecting how data workflows operate in enterprise settings. Gold-standard queries in this benchmark average close to 150 tokens, far surpassing previous datasets in both length and complexity. Notably, all SQLs are carefully annotated and reviewed for both semantic fidelity and execution correctness, ensuring challenges are both realistic and robust. As a result, Spider 2.0 highlighted the substantial gap between present-day LLMs’ capabilities and the demands of authentic, long-context database scenarios. In parallel, BIRD-CRITIC [^1] https://bird-critic.github.io/ [1] introduces a diagnostic dimension, comprising 600 tasks designed to assess the efficacy of LLMs in resolving real-world user issues across diverse SQL dialects. Together, these datasets push the boundaries of text-to-SQL research, emphasizing the challenges of complex, long-context SQL environments.
@@ III-A7 Specialized-domain dataset
Despite advancements in text-to-SQL research, there is still a shortage of specialized datasets tailored for domain-specific analysis. To address this gap, BULL [45] introduces a dedicated text-to-SQL benchmark for the financial sector. The dataset encompasses databases related to funds, stocks, and macroeconomic data, offering a comprehensive platform to evaluate and enhance text-to-SQL methods in financial applications, effectively tackling the financial domain’s unique text-to-SQL challenges.
@@ III-B2 Execution-based Metrics
    Then, $\textbf{R}(Y_{n},\hat{Y}_{n})=\sqrt{E(Y_{n})/E(\hat{Y}_{n})}$ denotes the relative execution efficiency of the predicted SQL in comparison to the ground-truth, where $E(\cdot)$ is the execution time of each SQL in the database. BIRD benchmark [1] ensures the stability of this metric by computing the average of $\textbf{R}(Y_{n},\hat{Y}_{n})$ over 100 runs for each example.
Most of the recent LLM-based text-to-SQL studies focus on BIRD and Spider-related datasets and execution-based metrics. The corresponding details can be found at Tab. III.
@@ IV-A In-context Learning
Through extensive and widely recognized research, ICL (also called prompt engineering) has been proven to play a decisive role in the performance of LLMs across various tasks [119, 26]. Consequently, it also impacts SQL generation across different prompt styles [8, 120]. The ICL implementation of the LLM-based text-to-SQL process to generate a predicted SQL query $\hat{Y}$, can be formulated as:
@@
where $k$ is the number of provided instances (examples) in few-shot prompting. $\mathcal{E}_{i}$ represents the $i$-th few-shot instance, each of which can be decomposed as $\mathcal{E}_{i}=\left(\mathcal{S}_{i},Q_{i},Y_{i}\right)$, where $Y_{i}$ is the gold SQL query of $Q_{i}$. The index $i\in\{1,\ldots,k\}$ also serves as the corresponding serial number for $\mathcal{S}_{i}$, $Q_{i}$, and $Y_{i}$. The OpenAI demonstration [^2] The prompt format follows the official document provided by the OpenAI platform: https://platform.openai.com/examples/default-sql-translate. provides an indicative example and regulation for the prompting process in LLM-based text-to-SQL [28]. The prompt designs and formats utilized by a substantial portion of the ICL methods discussed in this survey are developed based on the OpenAI regulation. Furthermore, most of these ICL methods incorporate zero-shot and few-shot prompting into their text-to-SQL frameworks.
@@ IV-A1 $\mathbf{C}_{0}$-Vanilla Prompting
TABLE II: A summary of representative methods used in the ICL paradigm for LLM-based text-to-SQL is provided. The complete table, along with a detailed categorization of existing well-designed methods denoted as $\mathbf{C}_{1:4}$, is available in Tab. III. *Note that SimpleDDL is a prompting style that provides only table and column names [120], rather than a tailored method.
@@
TABLE III: Well-designed methods in the ICL paradigm for LLM-based text-to-SQL are ordered by release time. The methods are grouped into four categories based on their implementation perspective: $\mathbf{C}_{1}$-Decomposition, $\mathbf{C}_{2}$-Prompt Optimization, $\mathbf{C}_{3}$-Reasoning Enhancement, and $\mathbf{C}_{4}$-Execution Refinement. Methods that fall into multiple categories will be introduced accordingly. The dataset column consists of six text-to-SQL datasets, which can be clarified by names: Spider [11], BIRD [1], Spider-DK [36] Spider-SYN [37], Spider-Realistic [38], and Spider 2.0 [42]
. We use the reference ID in the table due to column space limitations. *There are multiple applied LLMs in the corresponding method; the model with representative performance is presented. CoT method is introduced in multiple venues: NeurIPS’23 [1], EMNLP’23 [47], VLDB’24 [8], arXiv’24 [30]. Methods Applied LLMs Dataset Metrics $\mathbf{C}_{1}$ $\mathbf{C}_{2}$ $\mathbf{C}_{3}$ $\mathbf{C}_{4}$ Release Time Publication Venue MBR-Exec [78] Codex [11] EX Apr-2022 EMNLP’22 Coder-Reviewer [46] Codex [11] EX Nov-2022 ICML’23 LEVER [79] Codex [11] EX Feb-2023 ICML’23 Self-Debugging [80] StarCoder* [11] EX Apr-2023 ICLR’24 DESEM+P [63] ChatGPT [11, 37] EX Apr-2023 PRICAI’23 DIN-SQL [7] GPT-4* [1, 11, 42] EX, EM, VES Apr-2023 NeurIPS’23 CoT [47, 1, 8, 30] GPT-4 [11, 38, 1] EX, VES May-2023 Multiple Venues StructGPT [64] ChatGPT* [11, 37, 38] EX May-2023 EMNLP’23 SD+SA+Voting [65] ChatGPT* [11, 37, 38] EX May-2023 EMNLP’23 Findings QDecomp [47] Codex [11, 38] EX May-2023 EMNLP’23 Least-to-Most [47] Codex [11] EX May-2023 EMNLP’23 SQL-PaLM [74] PaLM-2 [11] EX May-2023 TMLR’24 RAG+SP&DRC [66] ChatGPT [11] EX Jul-2023 ICONIP’23 C3 [28] ChatGPT [11] EX Jul-2023 arXiv’23 DAIL-SQL [8] GPT-4* [11, 1, 38, 42] EX, EM, VES Aug-2023 VLDB’24 ODIS [67] Codex* [11] EX Oct-2023 EMNLP’23 Findings ACT-SQL [68] GPT-4* [11, 37] EX, EM Oct-2023 EMNLP’23 Findings MAC-SQL [48] GPT-4* [11, 1] EX, EM, VES Dec-2023 COLING’25 DEA-SQL [49] GPT-4 [11] EX Feb-2024 ACL’24 Findings FUSED [69] ChatGPT* [11] EX Feb-2024 EMNLP’24 Findings DELLM [70] GPT-4* [11, 1] EX, VES Feb-2024 ACL’24 Findings SGU-SQL [30] GPT-4* [11, 1] EX, EM Feb-2024 ICML’25 PoT [75] GPT-4* [11, 1] EX Feb-2024 arXiv’24 SQL-CRAFT [75] GPT-4* [11, 1] EX Feb-2024 arXiv’24 FUXI [76] GPT-4* [1] EX Feb-2024 EMNLP’24 MetaSQL [50] GPT-4* [11] EX, EM Feb-2024 ICDE’24 $\textit{{R}}^{3}$ [51] GPT-4* [11, 1] EX Feb-2024 ACL’25 Workshop PET-SQL [52] GPT-4 [11] EX Mar-2024 arXiv’24 PURPLE [53] GPT-4* [11, 37, 38] EX, EM Mar-2024 ICDE’24 Dubo-SQL [71] ChatGPT* [1] EX Apr-2024 arXiv’24 TA-SQL [54] GPT-4* [11, 1, 36, 38] EX Apr-2024 ACL’24 Findings MCS-SQL [55] GPT-4 [11, 1] EX, VES May-2024 COLING’25 CHESS [56] GPT-4o* [11, 1, 42] EX, VES May-2024 arXiv’24 SuperSQL [57] GPT-4* [11, 1] EX, EM, VES Jun-2024 VLDB’24 Distillery [58] GPT-4o-Mini* [1] EX Aug-2024 NeurIP’24 Workshop E-SQL [81] GPT-4o* [11, 1] EX, VES Sep-2024 arXiv’24 CHASE-SQL [77] Gemini-1.5* [11, 1] EX, VES Oct-2024 ICLR’25 RSL-SQL [59] DeepSeek-V3* [11, 1] EX, VES Nov-2024 arXiv’24 Spider-Agent [42] o1-preview* [42] EX Nov-2024 ICLR’25 ICRL [72] ChatGPT [11] EX Jan-2025 COLING’25 Gen-SQL [60] LLaMA-3* [11, 1] EX, EM, VES Jan-2025 COLING’25 SAFE-SQL [73] GPT-4o-Mini* [11] EX, EM Feb-2025 arXiv’25 ReFoRCE [61] o1-preview* [42] EX Feb-2025 ICLR’25 Workshop LinkAlign [62] DeepSeek-R1* [1, 42] EX, EM Mar-2025 arXiv’25